A fire arm of mass 0.1 kg fires a bullet of mass 1g that gains a velocity of 10 m/sec. The velocity gain of the fire arm due to recoil is 0.1m/sec. By how much foes the momentum of the fire arm increase, measured in kg m/sec?
Answer #1

Solution:

initial momentum of fire arm = 0 kgm/s

final momentum of fire arm = m * v = 0.1 * 0.1 = 0.01 kgm/s

=> increase in momentum = 0.01 - 0

=> increase in momentum = 0.01 kg m/s

thanks